# A Causal Framework to Measure and Mitigate Non-binary Treatment Discrimination

### v1

@@ C.1 Datasets and their preprocessing
Remarks. We converted all categorical features using label encoding. In Table 7, we summarize our datasets by identifying treatment features, sensitive attributes, and outcomes. We note that the treatment decision features for all the lending datasets in our work are continuous features. Hence, we use the difference between treatment values as $\mathcal{F}$ and the expectation $\mathcal{A}$ over each dataset $\mathcal{D}$ when considering disparity measures in Definitions 1 and 2. Furthermore, from Figure 4, we note that treatment feature distributions show significant overlap across the two sensitive groups. This observation shows that our overlap assumption is a reasonable one and holds in real-world settings.
@@
C.2 Losses
  1. 1.
    Loss Given Default (LGD): We estimate the loss incurred by the bank in the event an applicant defaults. Quantifying the loss will require additional analysis of other variables such as the value of collateral, instalments payments etc [25]. There are a number of different ways to estimate $LGD$ however a common approach considers the exposure at default and the recovery rate that is:
@@ C.2 Losses
    In our estimation, we assume that EAD is equivalent to the credit amount loaned out. EAD is a crucial metric that estimates the total loss a bank is exposed to when a loan defaults. However, it is important to note that the calculation of EAD can be more complex, as it may involve factors beyond the simple loan amount, such as commitment details [58].
    Finally, we formalize the expected loss given default for the bank as:
    LGD $=\mathbb{E}_{\mathcal{D}^{\prime}}[(1-y)\times\text{amount}]$ (12)
    where $y$ represents the applicant’s repayment probability.
  2. 2.
    Expected Simple Interest (ESI):
    annuity=(amount + interest)total durations(months)=(amount + $\text{annuity}=\frac{(\text{amount + interest})}{\text{total durations(months)}}=\frac{(\text{amount + interest})}{12\times\text{duration}}$
    Using this the simple interest can be calculated as :
    ESI(Home $\textbf{ESI}\text{(Home Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}[\text{annuity}\times 12\times\text{duration}-\text{amount}]$ (13)
    ESI(German $\textbf{ESI}\text{(German Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}\left[y\times\text{amount}\times\frac{\text{rate}}{100}\times\frac{\text{time}}{12}\right]$ (14)
Appendix D Synthetic loan
D.1 Synthetic dataset generation
Taking inspiration from [29], we consider the 7-variable loan dataset modeling the credit dataset [22] with an added outcome variable $Y$. Similar to [29], the equations are non-linear with non-additive exogenous variables. The nodes, gender $G$ and age $A$ are root-nodes (non-actionable), loan amount $L$, education-level $E$, income $I$, and savings $S$ are actionable with equations shown below. Additionally, we added a link between the savings $S$ to the loan amount and thresholded the outcome variable $Y$ directly on $G$ using two different thresholding scenarios. The symbol $\sigma(\cdot)$ represents the sigmoid function in the following equations. Distribution N (a,b) denotes the Gaussian distribution where a is the mean and b is the variance of the distribution, $\sqrt{b}$ denoting the standard deviation. Finally, Gamma (a, b) denotes the Gamma distribution where a is the concentration/shape and b is the scale, 1/b being the rate.
Root nodes : (15)
$f_{G}:G=U_{G},\quad U_{G}~{}\sim\operatorname{Bernoulli}(0.5)$
$f_{A}:A=-35+U_{A},\quad U_{A}~{}\sim\operatorname{Gamma}(10,3.5)$
Covariates:
$f_{E}:E=-0.5+\sigma(-1+0.5G+\sigma(0.1A)+U_{E}),$
$\quad\quad U_{E}\sim\mathcal{N}(0,0.25)$
$f_{I}:I=-4+0.1(A+35)+2G+GE+U_{I},U_{I}~{}\sim~{}\mathcal{N}(0,4)$
$f_{S}:S=-4+1.5\mathbb{I}\{I>0\}I+U_{S},U_{S}~{}\sim~{}\mathcal{N}(0,5)$
Treatments:
$f_{L}:L=1+0.01(A-5)(5-A)+2(1-G)+\beta S+U_{L},$
$\quad\quad U_{L}~{}\sim~{}\mathcal{N}(0,10),\beta\in\{0,0.03\}$
$f_{D}:D=-1+0.1A+3(1-G)+L+U_{D},U_{D}~{}\sim~{}\mathcal{N}(0,9)$
Outcome: two possibilities
$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS))$
$\geq 0.5+\gamma(1-G)\}\quad\textbf{OR}$
$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS)$
$+U_{Y}*\gamma(1-G))\geq 0.5\},U_{Y}~{}\sim~{}\mathcal{N}(0,5)$
[…]
$\alpha=1,\textbf{if}\left[\mathbb{I}\{I>0\}\land\mathbb{I}\{S>0\})\right]=1,$
$\alpha=-1,\textbf{otherwise}$
[…]
Table 8: Measuring treatment disparities and discrimination for synthetic data. Oracle uses perfect causal knowledge. Here, we have loan amount and duration as the treatment features. Results are reported as median values.
@@ D.1 Synthetic dataset generation
Balanced Female $\rightarrow$ Male TTD(-E) $-1.91$ $-4.44$ $-1.95$ $-4.94$ $0.19$ $0.00$ $0.18$ $0.00$
DTD(-E) $-1.62$ $-4.25$ $-2.00$ $-5.00$ $0.17$ $0.00$ $0.18$ $0.00$
Male $\rightarrow$ Female TTD(-E) $+2.14$ $+4.64$ $+1.95$ $+4.95$ $0.00$ $0.20$ $0.00$ $0.20$
DTD(-E) $+1.90$ $+4.05$ $+2.00$ $+5.00$ $0.00$ $0.17$ $0.00$ $0.20$
Imbalanced Female $\rightarrow$ Male TTD(-E) $-1.92$ $-4.61$ $-1.95$ $-4.95$ $0.09$ $0.00$ $0.09$ $0.00$
DTD(-E) $-1.55$ $-3.75$ $-2.00$ $-5.00$ $0.07$ $0.00$ $0.09$ $0.0$
Male $\rightarrow$ Female TTD(-E) $+2.00$ $+4.69$ $+1.95$ $+4.95$ $0.00$ $0.21$ $0.00$ $0.21$
DTD(-E) $+2.21$ $+5.45$ $+2.00$ $+5.00$ $0.00$ $0.24$ $0.00$ $0.21$
Table 9: Fairness in non-binary decision-making. Comparison of bank’s LGD and applicant’s ESI losses for balanced and unbalanced datasets under different policies.
Sensitive Group Balanced Dataset Unbalanced Dataset
LGD (DM) ESI (DM) LGD (DM) ESI (DM)

### v2

@@ B.3 Extending to Multiple Sensitive Attributes
B.4 Extended Discussion on Assumptions, Limitations, and Practical Considerations
We provide an extended discussion on the implicit assumptions of our existing causal framework and practical considerations for deployment.
Partial causal knowledge.
Our approach requires only partial knowledge of the underlying causal mechanisms—an assumption that is significantly weaker than that of many existing causal fairness methods (Kusner et al. 2017; Chiappa 2019), which typically assume full access to the causal graph. Specifically, we assume only the causal ordering among $S$, $X$, $Z$, and $Y$, as depicted in Figure 2. Both $X$ and $Z$ are treated as feature blocks, and we do not require knowledge of the internal causal ordering among covariates or treatment components. By leveraging CNFs to learn causal relations directly from observed data, our framework preserves theoretical guarantees even under partial causal knowledge, thereby improving its practical applicability.
No hidden confounders.
Consistent with prior work in causal inference (Khemakhem et al. 2021; Sánchez-Martin, Rateike, and Valera 2022; Javaloy, Sánchez-Martín, and Valera 2024) and causal fairness (Kusner et al. 2017; Chiappa 2019), we assume a data-generating process with no hidden confounders. From the decision-maker’s perspective, this corresponds to having complete access to the treatments $Z$, covariates $X$, and sensitive attributes $S$ used to assign $Z$, thereby removing confounding between $(Z,Y)$ and $(X,Z)$. In practice, however, decision-makers may lack full knowledge of all relevant covariates $X$ influencing $Y$, introducing hidden confounding between $X$ and $Y$. Such confounding does not affect our metrics for total disparities (Definitions 1 and 3) but can bias our direct disparity measures (Definitions 2 and 4), which rely on interventions on $X$.
External auditing and confounding.
Under an external auditor perspective, transparency in treatment assignment may be limited, and some components of $Z$ or the covariates $X$ used to determine $Z$ might be unobserved, introducing confounding among $X$, $Z$, and $Y$. Existing studies (Kilbertus et al. 2020a; Byun et al. 2024) provide fairness bounds under confounding, but they rely on strong assumptions or do not accommodate multivariate, non-binary treatments. Other approaches (Xu et al. 2021; Wang and Blei 2021; Miao et al. 2023) employ instrumental or proxy variables to mitigate hidden confounding, yet they are restricted to population-level effects and lack the individual-level specificity required for counterfactual analysis. This underscores the need for further research on fairness under confounding and the development of robust, individualized counterfactual methods. Incorporating recent techniques that enable reliable causal inference under specific types of confounding (Almodóvar et al. 2025) into our framework would be a promising direction for future work.
Overlap.
Beyond confounding, another key assumption underlying our framework and CNFs is overlap (Oberst et al. 2020). We assume that the intervened values in the SCM have non-zero support within the original data-generating process. This assumption generally holds in practice. For instance, our considered real-world lending datasets show substantial overlap in treatment distributions across sensitive groups. As seen in Figure 4, we have significant overlap between the distributions of treatment decisions for female and male applicants. Hence, assigning treatment decisions of the sensitive counterfactual is a realistic consideration for these datasets. Nevertheless, certain real-world contexts may exhibit lack of overlap between treatment distributions, for which extending our causal framework remains an important avenue for future research.
Counterfactual validity.
Accurate fairness evaluation depends on the fidelity of generated counterfactuals. Under valid causal assumptions, CNFs can produce highly accurate counterfactuals (Javaloy, Sánchez-Martín, and Valera 2024). However, validating these assumptions in real-world environments remains challenging. Our framework offers a practical advantage in this regard: since counterfactuals in our setting involve only treatment interventions — which are under the decision-maker’s control — they can be empirically evaluated via interventional studies. By deploying counterfactual treatment policies to specific subpopulations and measuring the resulting average effects, practitioners can assess and refine the quality of counterfactuals and the causal framework in practice.
Mitigation approaches.
Our mitigation strategy focuses on pre-processing to address treatment unfairness. Such techniques have proven effective in binary predictive fairness (Kamiran and Calders 2012) and provide a natural foundation for mitigating treatment disparities. However, pre-processing methods may not suit all scenarios and often lack theoretical guarantees. Future work should explore in-processing strategies capable of jointly optimizing fairness and utility objectives. Designing such approaches is non-trivial, as fair policy learning must account for binary decision fairness, non-binary treatment fairness, and stakeholder utility optimization. Achieving these objectives likely requires multi-objective optimization (Padh et al. 2021) for holistic fairness. Moreover, since treatment decisions and outcomes are typically conditioned on prior binary decisions (e.g., loan approvals), data may exhibit selection bias (Lakkaraju et al. 2017). Addressing this challenge may require online learning or policy adaptation methods (Kilbertus et al. 2020b; Rateike et al. 2022). Finally, future work should also assess whether in-processing methods can fully meet the multi-dimensional fairness requirements in decision-making (Favier and Calders 2025), or whether entirely new paradigms are needed.
B.5 Defining Stakeholder Losses
Here, we define the stakeholder loss functions we utilize to measure the utilities of decision-makers and decision-subjects for our mitigation strategy.
Loss Given Default (LGD):
We estimate the loss incurred by the bank in the event an applicant defaults. Quantifying the loss will require additional analysis of other variables, such as the value of collateral, instalments payments etc (Investopedia 2024). There are several different ways to estimate $LGD$; however, a common approach considers the exposure at default and the recovery rate which is:
LGD $=(1-\text{recovery rate})\times\text{exposure at default (EAD)}]$ (11)
In our estimation, we assume that EAD is equivalent to the credit amount loaned out. EAD is a crucial metric that estimates the total loss a bank is exposed to when a loan defaults. However, it is important to note that the calculation of EAD can be more complex, as it may involve factors beyond the simple loan amount, such as commitment details (Risk 2012).
Finally, we formalize the expected loss given default for the bank as:
LGD $=\mathbb{E}_{\mathcal{D}^{\prime}}[(1-y)\times\text{amount}]$ (12)
where $y$ represents the applicant’s repayment probability.
Expected Simple Interest (ESI):
$\text{annuity}=\frac{(\text{amount + interest})}{\text{total durations(months)}}=\frac{(\text{amount + interest})}{12\times\text{duration}}$
Using this the simple interest can be calculated as :
$\textbf{ESI}\text{(Home Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}[\text{annuity}\times 12\times\text{duration}-\text{amount}]$ (13)
$\textbf{ESI}\text{(German Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}\left[y\times\text{amount}\times\frac{\text{rate}}{100}\times\frac{\text{time}}{12}\right]$ (14)
B.6 Counterfactual Computation Pseudocodes
Generating sensitive counterfactual
Algorithm 1 computes the sensitive counterfactual following the causal relationships when ${\operatorname{do}}(S\rightarrow s^{CF})$. For this computation, we perform the single $\operatorname{do}$} operation on $S$ as ${\operatorname{do}}(S\rightarrow s^{CF})$.
Algorithm 1 compute_sensitive_counterfactual
Input: Factual data $\mathcal{D}^{F}$
Output: Sensitive counterfactual data $\mathcal{D}^{SCF}({\operatorname{do}}(S\rightarrow s^{CF}))$
1:Get sensitive feature values $s^{F}$ from $\mathcal{D}^{F}$.
3:Get data scaler from preparator.
[…]
5:return $\mathcal{D}^{SCF}$. $\triangleright$ clamp and floor categorical features before saving.
Computing downstream effect of counterfactual treatments
[…]
Algorithm 2 compute_path_specific_downstream_
effect_treatment_intervention
Input: Factual data $\mathcal{D}^{F}$, counterfactual data $\hat{\mathcal{D}}$
Parameters: Flows model, data preparator
Output: $)\mathcal{D}(s^{F},x^{F},{\operatorname{do}}({Z\rightarrow\hat{z}}$))
1:Get treatment indices $i_{\{Z\}}$ from preparator
2:Get values $\hat{z}$ from $\hat{\mathcal{D}}$ using $i_{\{Z\}}$
3:Get scaler from preparator
4:Clamp and floor $\hat{z}$ categorical features to proper value ranges
[…]
6:return $)\mathcal{D}(S^{F},X^{F},{\operatorname{do}}(Z\rightarrow\hat{z}$)) $\triangleright$ Extract $)Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow\hat{z}$)) and clamp, floor.
Computing direct effect of sensitive on the target
Algorithm 3 computes the direct effect $S$ on $Y$, i.e. ${\operatorname{do}}(S\rightarrow s^{CF})$ on the target $Y$ by computing $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$.
Algorithm 3 compute_path_specific_sensitive
_direct_label
Input: Factual data $\mathcal{D}^{F}$, sensitive counterfactual data $\mathcal{D}^{SCF}$ from Algorithm 1
Parameters: Flows model, data preparator
Output: $\mathcal{D}({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$
1:Get scaler from preparator
[…]
4:Get intervention indices $i_{\{X,T\}}$
5:Clamp, floor categorical dimensions in $\mathcal{D}^{F}[i_{\{X,T\}}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $\mathcal{D}^{F}$
6:Set data_tmp = clone($\mathcal{D}^{SCF}$)
[…]
9:Get combined $U_{\text{comb}}=[U^{SCF}_{S},U^{\mathcal{I}}_{\{X,Z\}},U^{F}_{Y}]$
[…]
11:return $\mathcal{D}({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ $\triangleright$ Extract $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ and clamp, floor.
Computing path-specific treatment
Algorithm 4 calculates the path-specific counterfactuals i.e., the path-specific treatment $z^{PS}=z({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$.
Algorithm 4 compute_path_specific_treatment
Input: Factual data $\mathcal{D}^{F}$, sensitive counterfactual data $\mathcal{D}^{SCF}$ from Algorithm 1
Parameters: Flows model, data preparator
Output: $\mathcal{D}^{PS}=D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$
1:Get scaler from preparator
[…]
4:Get intervention indices $i_{X}$
5:Clamp, floor categorical dimensions in $\mathcal{D}^{F}[i_{X}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $\mathcal{D}^{F}$
6:Set data_tmp = clone($\mathcal{D}^{SCF}$)
7:Set data_tmp[$i_{X}$] = intervention_vals[$i_{X}$]
[…]
11:return $\mathcal{D}^{PS}$ $\triangleright$ Extract $z^{PS}$, clamp and floor categorical features before saving.
Computing downstream effect of $z^{PS}$
Appendix C Experimental Setup
C.1 Datasets and Preprocessing
  • HMDA: This dataset regarding mortgage applications covers several states in the United States, but we focused on just two states (New York and Texas) in the year 2017. Following similar preprocessing steps as done in (Cooper et al. 2024), we binarized the gender column as $\{Male:1,Female:0\}$. Also, we binarized the race with originally $8$ possible values but utilized only $\{NativeWhite:0,White:1\}$. For the action taken with $8$ possible values, we filtered out 38{3-8}. We kept the outcome as a $1$ Loan originated to mean the applicant and the bank decision agreed, and $2$, which is the application approved but not accepted, to mean a disagreement. In dealing with missing values we removed columns with all NaNs and information regarding the co-applicants. After the preprocessing, we had a total dataset of size $399354$ with $17$ features.
  • Home Credit: We downloaded this dataset from the kaggle competition. We used the application training datasets and selected features for our study based on the feature importance as done in (Matthys 2019). We utilized the resulting features selected by their random forest classifier and information entropy models. Based on this, we selected the following features : “EXT SOURCE 2, EXT SOURCE 3, EXT SOURCE 1, DAYS BIRTH, NAME EDUCATION TYPE, CODE GENDER, DAYS EMPLOYED, NAME INCOME TYPE, ORGANIZATION TYPE, AMT CREDIT, AMT ANNUITY, AMT GOODS PRICE, REGION POPULATION RELATIVE, TARGET”. Based on this subset of data, we dropped the missing values in the dataset and converted the days of birth and days of employment to years using $\frac{\text{days of birth/employment}}{-365}$. We divided it by 365-365 because these values were negative. The complete dataset has a size of $98859$ with $14$ features.
@@ C.1 Datasets and Preprocessing
Remarks. We converted all categorical features using label encoding. In Table 5, we summarize our datasets by identifying treatment features, sensitive attributes, and outcomes. We note that the treatment decision features for all the lending datasets in our work are continuous features. Hence, we use the difference between treatment values as $\mathcal{F}$ and the expectation $\mathcal{A}$ over each dataset $\mathcal{D}$ when considering disparity measures in Definitions 1 and 2.
@@
C.2 Hyper-Parameter Tuning and Training of Causal Normalising Flows
For the causal normalizing flows (Javaloy, Sánchez-Martín, and Valera 2024), we use the original work as inspiration and tune for the flow models’ inner dimensionality and learning rate Table 6. All other parameters are fixed, which are as follows: Adam optimizer with plateau scheduler, activation as ELU, layer type was Neural Spline Flow (NSF) and all the experiments were run for a maximum of $1000$ epochs. For other parameters, we refer to the GitHub code  [^2] https://github.com/psanch21/causal-flows. The number of nodes is specified in Table 7. An example of causal graph input to causal normalizing flows is given in Figure 5.
Dataset dim-inner lr log-prob
HMDA-NY $32\times 32$ $10^{-2}$ -61.67
HMDA-TX $32\times 32$ $10^{-2}$ -63.49
Home Credit $64$ $10^{-2}$ -41.87
German Credit $64$ $10^{-3}$ -14.68
Synthetic Loan $16\times 16\times 16\times 16$ $10^{-3}$ -16.51
Dataset #rootnodes #covariates #treatment
HMDA-NY 2 12 2
HMDA-TX 2 12 2
Home Credit 2 9 2
German Credit 2 15 3
Synthetic Loan 2 3 2
C.3 Hardware
All experiments of hyper-parameter tuning and training of causal normalizing flows are done on an Ubuntu workstation with an Intel CPU. The code primarily uses PyTorch CPU and Python. We set the seed for PyTorch and Python to a fixed number and ran all experiments. A fixed seed is also used to split datasets into train, validation, and test sets - 80%, 10%, 10%. This seed is fed as input in a parameter file to train and run evaluations.
AG
EISLDYSensitive SSOutcome YYBlock XXBlock $Z$
Appendix D Synthetic Validation Evaluation
In this section, we consider a synthetic data generation process where we have access to the ground-truth generative process. Using this dataset, we aim to validate the performance of our causal framework, comparing it to the ground-truth oracle that has access to the exact causal relations. This analysis shows that, considering the underlying assumptions hold, our causal framework leveraging the CNF can capture unknown causal relations and reliably help audit existing treatment disparities and also mitigate potential discriminatory effects of such treatment decision disparities.
D.1 Synthetic Dataset Generation
Taking inspiration from (Karimi et al. 2020; Majumdar and Valera 2024), we consider the 7-variable loan dataset modeling the credit dataset (Hofmann 1994) with an added outcome variable $Y$ (Equation 6). Similar to (Karimi et al. 2020; Majumdar and Valera 2024), the equations are non-linear with non-additive exogenous variables. Gender $G$ and age $A$ are root-nodes, loan amount $L$, education-level $E$, income $I$, and savings $S$ are actionable with equations shown below. Additionally, we added a link between the savings $S$ to the loan amount and thresholded the outcome variable $Y$ directly on $G$ using two different thresholding scenarios. $\sigma(\cdot)$ represents the sigmoid function in the following equations. $\mathcal{N}(a,b)$ denotes the Gaussian distribution where a is the mean and b is the variance of the distribution, $\sqrt{b}$ denoting the standard deviation. Finally, $\operatorname{Gamma}(a,b)$ denotes the Gamma distribution where a is the concentration/shape and b is the scale, 1/b being the rate.
@@ D.1 Synthetic Dataset Generation
Figure 6: Synthetic data generating process.
@@
Balanced Female $\rightarrow$ Male TTD(-E) 1.91-1.91 4.44-4.44 1.95-1.95 4.94-4.94 $0.19$ $0.00$ $0.18$ $0.00$
DTD(-E) 1.62-1.62 4.25-4.25 2.00-2.00 5.00-5.00 $0.17$ $0.00$ $0.18$ $0.00$
Male $\rightarrow$ Female TTD(-E) +2.14+2.14 +4.64+4.64 +1.95+1.95 +4.95+4.95 $0.00$ $0.20$ $0.00$ $0.20$
DTD(-E) +1.90+1.90 +4.05+4.05 +2.00+2.00 +5.00+5.00 $0.00$ $0.17$ $0.00$ $0.20$
Imbalanced Female $\rightarrow$ Male TTD(-E) 1.92-1.92 4.61-4.61 1.95-1.95 4.95-4.95 $0.09$ $0.00$ $0.09$ $0.00$
DTD(-E) 1.55-1.55 3.75-3.75 2.00-2.00 5.00-5.00 $0.07$ $0.00$ $0.09$ $0.0$
Male $\rightarrow$ Female TTD(-E) +2.00+2.00 +4.69+4.69 +1.95+1.95 +4.95+4.95 $0.00$ $0.21$ $0.00$ $0.21$
DTD(-E) +2.21+2.21 +5.45+5.45 +2.00+2.00 +5.00+5.00 $0.00$ $0.24$ $0.00$ $0.21$
Table 8: Measuring treatment disparities and discrimination for synthetic data. Oracle uses perfect causal knowledge. Here, we have loan amount and duration as the treatment features. Results are reported as median values. Our framework using CNF can accurately capture disparities without access to ground-truth causal mechanisms.
Sens. Group Balanced Data Unbalanced Data
LGD ESI LGD ESI
@@
Table 9: Fairness in non-binary decision-making. Comparison of the bank’s LGD and the applicant’s ESI losses (in DM) for balanced and unbalanced datasets under different treatments.